OpenStudy (anonymous):

4 cos^2 x =2+2sinx .. wht is the solution of the equation in the interval [0,2pie) ? i got no solution but i m nt sure

OpenStudy (anonymous):

\[ 4 \cos^2 (x) =2+2\sin (x)\\ 4( 1- \sin^2(x))= 2+2\sin (x)\\ 4 -4\sin^2(x)= 2+2\sin (x)=0\\ -4 +4\sin^2(x)+ 2+2\sin (x)=0\\ 4\sin^2(x)+2\sin (x)-2=0\\ 2\sin^2(x)+\sin (x)-1=0\\ \] Can you finish it now?

OpenStudy (anonymous):

Use the quadratic formula to find \[ \sin(x) = -1\\ \sin(x) = \frac 1 2 \]

OpenStudy (anonymous):

\[ x= \frac {3 \pi}2\\ x= \frac \pi 6\\ x= \frac {5\pi}6 \]

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

hey eliasaab if u could just check my answers for this question it would be a great help

OpenStudy (anonymous):

the sign of the tan should be negative the sign of the sec should be positive

OpenStudy (anonymous):

You can practice trig problems on my site http://www.math.missouri.edu/escgi/mucgi-bin/munew.cgi?variable=trig